# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

### v1

@@ 5.1 Synthetic data generation
To distill knowledge from the proposed model, we collected the outputs of over a hundred NLG systems and applied OpeNLGaugeens to produce synthetic evaluation data. We include NLG outputs produced on 15 datasets covering five task categories and almost 40 aspects. Table 2 shows the basic statistics of the constructed dataset.
The NLG outputs included cover the following tasks: data-to-text, summarization, question answering, dialogue response generation and story generation. These were chosen to represent a diverse range of tasks, each with unique objectives, input-output relationships and evaluation aspects. Data-to-text represents tasks where input is transformed from one representation to another. Summarization compresses information from a longer text into a shorter version. Question answering extracts relevant information from a context in response to a specific query. Finally, dialogue response generation and story generation create outputs with novel content. Note that the underlying datasets are used only as input for NLG systems and do not include human evaluations of any kind.
For each task, we select a set of relevant evaluation aspects with their definitions. We consider two aspects distinct if they are associated with different tasks, even if they share the same name. For instance, coherence in dialogue response generation refers to the coherence of the response with the dialogue history, while in summarization, it refers to the internal coherence of a summary itself. We list aspects and datasets for each task in Appendix A.
To obtain outputs with varying quality and diverse types of errors, we sample outputs from a variety of systems, ranging from rule-based approaches to the latest state-of-the-art LLMs. For older systems, we use pre-generated outputs from existing datasets, while outputs from more recent systems including LLMs are newly generated. An overview of the evaluated systems is shown in Appendix B.
To limit computational requirements for dataset generation, we decided to apply sampling that ensures data diversity and broad coverage of different NLG systems and aspects while significantly reducing the size of the dataset. For each input, we randomly sample the outputs of $N$ systems, and then $M$ aspects are sampled for each input-output pair. This results in $N\,\times\,M$ (input, output, aspect) triples for each input, which are then annotated by OpeNLGaugeens to obtain synthetic annotations. For most tasks, we set $N=4$ and $M$ = 3. This sampling strategy aims for a balanced distribution of inputs, NLG system outputs and evaluation aspects. Additionally, it ensures exposure to different outputs for the same input, i.e., it should not prime models to evaluate based solely on patterns in the inputs. Finally, by presenting multiple aspects for the same input-output pair, models are encouraged to learn differences in output quality across various evaluation aspects.
@@
Table 2: Dataset statistics: number of source datasets (Src.), systems (Sys.), evaluation aspects (Asp.) and training examples for each task.
@@ 5.2 Fine-tuning procedure
As the model is expected to learn the task from training data, we used a special prompt template for fine-tuning. The prompt contains only a brief description of the task, the definition of the evaluated aspect, as well as the input and output to be evaluated. The prompt is provided in the Appendix E.
@@ 6 Experimental Setup
6.1 Models
Ensemble of LLMs
The ensemble consists of six open-weight LLMs: five annotators and one supervisor. Each selected LLM is distributed under a license that permits at least non-commercial use and allows the model’s outputs to be used as training data. At the time of performing the experiments, these models ranked among the top-performing open-weight LLMs on the Chatbot Arena Leaderboard [^4] https://lmarena.ai/?leaderboard (Chiang et al., 2024).
The annotator models include the following: Llama 3.1 Nemotron 70B (Wang et al., 2024), Qwen 2.5 70B (Yang et al., 2024), Gemma 2 27B (Gemma Team et al., 2024), Command R+ 104B (Cohere For AI, 2024), and Mistral Large 2 123B [^5] https://mistral.ai/news/mistral-large-2407/. We apply Llama 3.3 70B (Dubey et al., 2024) as the supervisor model. To address computational constraints, we use quantized versions available through the Ollama platform. [^6] https://ollama.com/ For synthetic data generation, we set the temperature to zero to obtain consistent results. For details, see Appendix C.
Fine-tuning parameters
We use LoRA with rank 16 and alpha 32 to fine-tune Llama 3.1 8B. The model is trained for one epoch with a learning rate of 2e-4, using AdamW optimizer (Loshchilov and Hutter, 2017) with a weight decay of 0.01. We apply a linear learning rate schedule with a warm-up period corresponding to the first 5% of training steps. The batch size is set to 16. This setup enables resource-efficient training that requires only around 20GB of VRAM, completing one training epoch in just six hours on a single A40 GPU.
@@ 6.2 Evaluation Datasets
We used seven popular meta-evaluation datasets to assess the correlation of our approach with human judgements (see their detailed description in Appendix D). The datasets cover the following tasks: summarization – SummEval (Fabbri et al., 2021b), QAGS (Wang et al., 2020a); story generation – HANNA (Chhun et al., 2022); data-to-text – SFRES and SFHOT (Wen et al., 2015), text simplification – Wiki-DA (Alva-Manchego et al., 2021) and dialogue response generation – TopicalChat (Gopalakrishnan et al., 2023). Note that text simplification is a task unseen during training and TopicalChat contains an unseen aspect (groundedness). For human evaluation of error spans, we used the RotoWire dataset (Thomson and Reiter, 2020) containing annotations from a data-to-text task in the basketball domain, which was unseen during training by the evaluated metrics.
System Error span Error No span given Not an error Hallucination Total
Explanation
OpeNLGaugeens Correct 218 9 3 0 230
Partially correct 41 0 0 1 42
Incomplete 20 0 1 0 21
Vague 4 2 1 0 7
Incorrect 20 3 32 1 56
Not an error 0 0 10 0 10
Total 303 14 47 2 366
Span OK (%) 83
Exp. correct (%) 63
OpeNLGaugeft Correct 108 3 4 4 119
Partially correct 60 1 2 0 63
Incomplete 9 0 0 0 9
Vague 3 9 0 0 12
Incorrect 41 7 42 6 96
Not an error 0 1 8 0 9
Total 221 21 56 10 308
Span OK (%) 72
Exp. correct (%) 39
TigerScore Correct 8 17 0 0 25
Partially correct 5 2 0 0 7
Incomplete 2 2 0 0 4
Vague 0 3 0 0 3
Incorrect 5 16 2 0 23
Not an error 1 2 0 0 3
Total 21 42 2 0 65
Span OK (%) 32
Exp. correct (%) 38
Table 4: Results of human evaluation of error-span explanations. Each table shows absolute occurrence counts of different error span and explanation validity (see Section 7.2), with overall proportions of correct annotation given below.
@@ 6.3 Baselines
We compare our methods to a variety of commonly used evaluation metrics, including traditional metrics like ROUGE (Lin, 2004b), BLEU (Papineni et al., 2002b) and METEOR (Agarwal and Lavie, 2008), distance-based metrics like MoverScore (Zhao et al., 2019), BERTScore (Zhang et al., 2019), and trained metrics like BARTScore (Yuan et al., 2021) and UniEval (Zhong et al., 2022). We also include LLM-based metrics like GPTScore (Fu et al., 2024), G-Eval (Liu et al., 2023), Prometheus (Kim et al., 2024b), Auto-J (Li et al., 2023a), InstructScore (Xu et al., 2023b), TIGERScore (Jiang et al., 2023) and Themis (Hu et al., 2024b). Additionally, we report the performance of LLMs directly prompted to evaluate the outputs: GPT-3.5 and GPT-4 (see Chiang and Lee (2023b)). To measure the improvement of OpeNLGaugeft relative to the base model, we also include Llama 3.1 8B as an additional baseline. The specific metrics reported differ depending on the evaluated dataset and NLG task.
Additionally, our comparisons include some task-specific and aspect-specific metrics: QAGS (Wang et al., 2020a) and FactCC (Kryscinski et al., 2020) for evaluating factual consistency in summarization; SARI (Xu et al., 2016) and LENS (Maddela et al., 2022) for text simplification, and USR (Mehri and Eskenazi, 2020b) for dialogue response generation tasks. The latter metric has several variants; we use the variant with the best Pearson correlation for each aspect in our comparison.
@@ 7.1 Score Correlation with Humans
Figure 3: Results of ablation experiments on QAGS and TopicalChat for OpeNLGaugeens. The plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the ensemble with the original prompt and the corresponding ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average for all aspects.
The results for factual consistency on QAGS are presented in Table 3 and the results for other datasets are available in Appendix F. On QAGS, OpeNLGaugeens achieves the highest average performance on both Pearson and Spearman correlations, while on Kendall it is outperformed by Themis. The distilled version of our metric was consistently the third best measure. Notably, it outperformed the proprietary GPT-4 on this task.
On the other summarization benchmark, SummEval, the best performing metric was Themis, followed by OpeNLGaugeens. Comparing our approach to another method that provides a similar level of explainability (error span annotations), TigerScore, we observe 15 p.p. improvements on the Pearson correlation averaged over all aspects. The smaller OpeNLGaugeft outperformed all other fine-tuned LLM-based metrics except Themis, and also outperformed GPT-3.5 and the version of G-Eval that used it by a large margin.
On data-to-text problems, OpeNLGaugeens excelled in evaluating naturalness of generated texts, while OpeNLGaugeft showed as the strongest evaluator of informativeness on the SFRES dataset. Interestingly, on data-to-text problems, the distilled version of our model achieved slightly better results on average than our ensemble. A similar situation is observed on the story generation dataset, where the distilled version obtained better average scores on Spearman and Kendal correlations, and closely followed OpeNLGaugeens on Pearson.
@@
On TopicalChat, the proprietary GPT-4 emerged as the strongest evaluator, with our OpeNLGaugeens performing between that model and GPT-3.5. OpeNLGaugeft achieved almost 25 p.p. improvement in both Pearson and Spearman correlations compared to its Llama 3.1 backbone. Most importantly, a large improvement is also observed on the aspect not seen during training (groudedness), where the model outperformed all fine-tuned LLM-based metrics except Themis on Pearson correlation.
Generalization to unseen tasks
For text simplification, our ensemble shows superior performance across the board, outperforming LENS, a strong baseline metric specialized for this task. OpeNLGaugeft, although not originally trained on the task, outperforms all baseline metrics on average correlation and achieves a particularly high score on meaning preservation.
@@ 7.2 Human Evaluation of Error Spans
We performed a small in-house human evaluation study to compare the quality of explanations obtained by OpeNLGauge and TigerScore, another LLM-based metric that also provides error-span annotations (see Table 1). For this purpose, we used a data-to-text task in the basketball domain (RotoWire dataset (Thomson and Reiter, 2020)). Five expert annotators evaluated the output of all three systems on 40 instances (a total of 120 outputs). We asked the annotators to: (1) evaluate provided error-spans, marking them as correct error spans, spans not containing an error, hallucinated spans, and situations where no span was provided; (2) evaluate generated explanations, marking them as correct, partially correct, incomplete, vague, incorrect, or texts that do not describe an error.
The results are shown in Table 4. Both OpeNLGaugeft and OpeNLGaugeens were over two times more accurate than TigerScore at annotating error spans, while finding over ten times more correct error spans. The task of providing accurate error explanations was more difficult for all the approaches evaluated. OpeNLGaugeens achieved the highest performance and was almost twice as accurate as TigerScore and OpeNLGaugeft, which achieved similar accuracies.
@@ 7.3 Ablation Experiments
To better analyze the proposed approach, we perform ablation experiments on the selected prompt and ensemble structure.
@@
We explore the effect of using different scales for the overall score and error severity scales: integer scales for both (Ablation 1), integer scale for overall and categorical scale for severity (Ablation 2), categorical scale for both (Ablation 3). Recall that OpeNLGauge uses a categorical scale for overall score and integer scale for severity – see Section 4.
The differences in the correlations between the full prompt and the ablations are summarized in Figure 3, for detailed results see Appendix G.1. Overall, the change of scale has little effect on the average correlation of the whole ensemble, but it has a dramatic effect on some individual LLMs and aspects. This illustrates the ability of the ensemble to compensate for weaknesses in individual models.
@@
We also analyze the effect of ensemble size and the influence of its particular components on the correlations with human scores by recomputing the results for all ensemble combinations. The results are presented in Appendix G.2. For Wiki-DA, performance increases with ensemble size, with the full ensemble being the best combination. For other datasets, there are a few smaller combinations that actually rank higher, but none is consistently better than the full ensemble.

### v3

@@ 5.1 Synthetic data generation
To distill knowledge from the ensemble, we collected the outputs of over a hundred NLG systems and applied OpeNLGaugeens to produce synthetic evaluation data. We include NLG outputs produced on 15 datasets covering five task categories and almost 40 aspects. Table 2 shows the basic statistics of the constructed dataset.
@@
Table 2: Training dataset statistics for OpeNLGaugeft: number of source datasets (Src.), systems (Sys.), evaluation aspects (Asp.) and training examples for each task.
The NLG outputs included cover the following tasks: data-to-text, summarization, question answering, dialogue response generation and story generation. These were chosen to represent a diverse range of tasks, each with unique objectives, input-output relationships and evaluation aspects. Note that the underlying datasets are used only as inputs for NLG systems and do not include human evaluations of any kind.
For each task, we select a set of relevant evaluation aspects with their definitions. We consider two aspects distinct if they are associated with different tasks. For instance, coherence in dialogue refers to coherence of the response with respect to the dialogue history, while in summarization it refers to the internal coherence of a summary. We list the datasets and aspects for each task in Appendix A.
To obtain output texts with varying quality and diverse types of errors, we sample outputs from a variety of systems, ranging from rule-based approaches to state-of-the-art LLMs. For older systems, we use pre-generated outputs from existing datasets, while outputs from more recent systems including LLMs are newly generated. An overview of the evaluated systems is shown in Appendix B.
To limit computational requirements for dataset generation, we apply a sampling procedure that ensures data diversity and broad coverage of different NLG systems and aspects while significantly reducing dataset size. For each input, we randomly sample the outputs of $N$ systems, followed by sampling $M$ aspects for each input-output pair. The sampling includes all aspects described in Appendix A and all systems listed in Appendix B. This results in $N\,\times\,M$ (input, output, aspect) triples for each input, which are then passed to OpeNLGaugeens to obtain synthetic annotations. For most tasks, we set $N=4$ and $M$ = 3. This sampling strategy aims for a balanced distribution of inputs, NLG system outputs and evaluation aspects. It also ensures exposure to different outputs for the same input, i.e., it should not prime models to evaluate based solely on patterns in the inputs. Finally, by presenting multiple aspects for the same input-output pair, models are encouraged to learn differences in output quality between different evaluation aspects.
To keep the merged evaluation outputs internally consistent, we remove outliers before merging, as described in Section 4. Table 12 in Appendix D shows the proportions of outliers detected for all LLMs and task categories (3.4% on average).
@@ 5.2 Fine-tuning Procedure
As the model is expected to learn the task from training data, we used a simpler prompt template for fine-tuning, with a brief description of the task, the definition of the evaluated aspect, and the input and output to be evaluated (see Appendix F).
@@ 6 Experimental Setup
6.1 Ensemble
The ensemble consists of six open-weight LLMs: five annotators and one consolidator. Each selected LLM is distributed under a license that permits at least non-commercial use and allows the model’s outputs to be used as training data. At the time of the experiments, these models ranked among the top-performing open-weight LLMs on the Chatbot Arena Leaderboard [^4] https://lmarena.ai/leaderboard (chiang2024chatbot).
The annotator models include the following: Llama 3.1 Nemotron 70B (wang2024helpsteer2), Qwen 2.5 72B (yang2024qwen2), Gemma 2 27B (gemma2024gemma), Command R+ 104B (cohere_for_ai_2024), and Mistral Large 2 123B [^5] https://mistral.ai/news/mistral-large-2407/. We apply Llama 3.3 70B (dubey2024llama) as the consolidator model. To address computational constraints, we use quantized versions available through the Ollama platform. [^6] https://ollama.com/ For synthetic data generation, we set the temperature to zero to obtain consistent results. For details on the models, see Appendix C.
6.2 Distillation
To produce OpeNLGaugeft, we apply LoRA hu2022lora with rank 16 and alpha 32 to fine-tune the instruction-tuned version of Llama 3.1 8B. The model is trained for one epoch with a learning rate of 2e-4, using AdamW optimizer (loshchilov2017weight) with a weight decay of 0.01. We apply a linear learning rate schedule with a warm-up period corresponding to the first 5% of training steps. The batch size is set to 16. This setup enables resource-efficient training that requires only around 20GB of VRAM, completing one training epoch in just six hours on a single A40 GPU.
@@ 6.3 Evaluation Datasets
We used seven popular meta-evaluation datasets to assess how our metric correlates with human judgments (for detailed descriptions of these datasets, including details on aggregation of human scores, see Appendix E). The datasets cover the following tasks: summarization – SummEval (fabbri2021summeval), QAGS (wang2020asking); story generation – HANNA (chhun2022human); data-to-text – SFRES and SFHOT (wen2015semantically), text simplification – Wiki-DA (alva2021suitability) and dialogue generation – TopicalChat (gopalakrishnan2023topical). For OpeNLGaugeft, text simplification is a task unseen during training and TopicalChat contains an unseen aspect (groundedness). For human evaluation of error spans, we used the RotoWire dataset (thomson2020gold) with annotations from a data-to-text task in the basketball domain, a task unseen during training by the evaluated fine-tuned metrics.
Dataset Llama 3.1 OpeNLGaugeft $\Delta$
QAGS 0.236 0.651 +0.415
SummEval 0.186 0.502 +0.316
TopicalChat 0.309 0.578 +0.269
SFRES/SFHOT 0.108 0.315 +0.207
HANNA 0.150 0.425 +0.275
Wiki-DA 0.405 0.789 +0.384
Table 4: Comparison of Spearman ($\rho$) correlations of the backbone model (Llama 3.1 8B) and our metric OpeNLGaugeft fine-tuned from the backbone on the dataset described in Section 5.1. For each dataset, the correlations are averaged across all evaluated aspects.
@@ 6.4 Baselines
We compare our methods with a variety of commonly used evaluation metrics, including traditional metrics such as ROUGE (lin-2004-rouge), BLEU (papineni2002bleu) and METEOR (agarwal2008meteor), distance-based metrics MoverScore (zhao-etal-2019-moverscore) and BERTScore (zhang2019bertscore), and trained metrics BARTScore (yuan2021bartscore) and UniEval (zhong2022towards). We also include the following LLM-based metrics: GPTScore (fu2024gptscore), G-Eval (liu-etal-2023-g), LLM Evaluation (chiang-lee-2023-closer), Prometheus (kim2023prometheus), Auto-J (li2023generative), InstructScore (xu-etal-2023-instructscore), TIGERScore (jiang2023tigerscore) and Themis (hu2024themis). To measure the improvement of OpeNLGaugeft relative to the base model, we include the instruction-tuned version of Llama 3.1 8B as an additional baseline. The specific metrics reported differ depending on the dataset and the evaluated NLG task.
Additionally, our comparisons include some task-specific and aspect-specific metrics: QAGS (wang2020asking) and FactCC (kryscinski-etal-2020-evaluating) for evaluating factual consistency in summarization, SARI (xu-etal-2016-optimizing) and LENS (maddela2022lens) for text simplification, and USR (mehri2020usr) for dialogue response generation tasks. The latter metric has several variants; we use the variant with the best Pearson correlation for each aspect in our comparison.
@@ 7 Results
Figure 3: Results of human evaluation of error spans and explanations. Top half of each bar: Error spans marked as correct or incorrect (hallucinated spans, no span provided, or spans without errors). Bottom half: Explanations marked as correct, partial (partially correct or incomplete) or incorrect (not addressing actual errors, vague or incorrect). The differences between TigerScore and OpeNLGaugeens are statistically significant (t-test, $p<0.05$). See Table 13 for more details.
Figure 4: Ablation results on QAGS and TopicalChat for OpeNLGaugeens. Plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the ensemble with the original prompt and the corresponding ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average for all aspects.
@@ 7.1 Score Correlation with Humans
The results for factual consistency on QAGS are presented in Table 3 and the results for other datasets are available in Tables 1420 in Appendix 13. On QAGS, OpeNLGaugeens achieves the highest average performance on both Pearson’s $r$ and Spearman’s $\rho$. On Kendall’s $\tau$, our method is outperformed by Themis, which can be attributed to different score granularities used by these two methods (see Section G.1 for a discussion). The distilled version of our metric was consistently the third best measure. Notably, it outperformed the metrics based on the proprietary GPT-4 on this task.
On SummEval (Table 14), the best performing metric was Themis, closely followed by OpeNLGaugeens. However, note that training data for Themis include almost 62,000 human-annotated examples for the summarization task. Comparing our approach to TigerScore, another method that provides a similar level of explainability (error span annotations), we observe 15 p.p. improvements on Spearman’s $\rho$ averaged over all aspects. The smaller OpeNLGaugeft outperformed all other fine-tuned LLM-based metrics except Themis, and also surpassed metrics based on prompting GPT-3.5 by a large margin.
On TopicalChat (Table 15), the LLM Evaluation metric based on GPT-4 emerged as the strongest evaluator, while OpeNLGaugeens ranked between this method and its GPT-3.5-based version. OpeNLGaugeft achieved 27 p.p. improvement in Spearman’s $\rho$ compared to its Llama 3.1 backbone.
In data-to-text, OpeNLGaugeens excelled in evaluating naturalness, while OpeNLGaugeft stands out as the strongest evaluator of informativeness on the SFRES dataset (Table 16). Interestingly, on data-to-text problems, the distilled version of our metric achieved slightly better results on average than our ensemble. A similar situation is observed for story generation (Tables 1719), where the distilled version obtained better average scores on Spearman’s $\rho$ and Kendall’s $\tau$, and closely followed OpeNLGaugeens on Pearson’s $r$.
In text simplification (Table 20), our ensemble achieved superior performance across the board, outperforming LENS, a strong baseline metric specialized for this task.
Additionally, Table 4 presents a comparison of OpeNLGaugeft with Llama 3.1 8B (instruct), indicating a considerable improvement over the prompted model.
Generalization to unseen tasks
Although not originally trained on text simplification, OpeNLGaugeft outperforms all baseline metrics and individual LLMs in averaged Pearson correlation on the Wiki-DA dataset (Table 20) and achieves a particularly high score on meaning preservation.
@@
On TopicalChat (Table 15), a large improvement for OpeNLGaugeft over Llama 3.1 8B is observed on groundedness, which is an aspect unseen during training. OpeNLGaugeft also surpassed most fine-tuned metrics and most individual LLMs on this aspect. Moreover, Wiki-DA contains additional unseen aspects (meaning preservation and simplicity) on which OpeNLGaugeft shows considerable improvement over Llama 3.1 8B and outperforms most of the individual larger LLMs.
@@ 7.2 Human Evaluation of Error Spans
We performed a small in-house human evaluation study to compare the quality of explanations obtained by OpeNLGauge and TigerScore, another LLM-based metric that also provides error-span annotations (see Table 1). For this purpose, we used a data-to-text task in the basketball domain (thomson2020gold). Five expert annotators evaluated the output of all three systems on 40 instances (a total of 120 outputs and 950 error spans). We asked the annotators to: (1) evaluate provided error spans, marking them as correctly identified, not containing an error, hallucinated, and situations where no span was provided; (2) evaluate generated explanations, marking them as correct, partially correct, incomplete, vague, incorrect, or texts that do not describe an error.
To assess reliability of our human annotation, we computed Cohen’s $\kappa$ coefficient (cohen1960coefficient) of inter-annotator agreement on 50 error spans with double annotations. We obtained $\kappa=0.82$ for the evaluation of error spans, and $\kappa=0.46$ for error explanations.
The results are shown in Figure 3, with more details provided in Table 13 in Appendix 13. Both OpeNLGaugeft and OpeNLGaugeens are over twice more accurate than TigerScore at annotating error spans, while finding over ten times more correct error spans. The task of providing accurate error explanations was more difficult for all the approaches evaluated. OpeNLGaugeens achieved the highest performance and was almost twice as accurate as TigerScore and OpeNLGaugeft, which achieved similar accuracies.
@@ 7.3 Ablation Experiments
We explore the effect of using different scales for the overall score and error severity: integer scales for both (Ablation 1), integer scale for overall score and categorical scale for severity (Ablation 2), and categorical scale for both (Ablation 3). Recall that OpeNLGauge uses a categorical scale for overall score and integer scale for error severity – see Section 4.
Correlation differences between the full prompt and the ablations are summarized in Figure 4, for detailed results see Appendix H.1. Overall, the change of scale has little effect on the average correlation of the whole ensemble, but it has a dramatic effect on some individual LLMs and aspects. This illustrates the ability of the ensemble to compensate for weaknesses in individual annotator models.
@@
We also analyze the effect of ensemble size and the influence of its particular components on the correlations with human scores by recomputing the results for all ensemble combinations. The results are presented in Appendix H.2. For Wiki-DA, performance increases with ensemble size, with the full ensemble being the best combination. For other datasets, there are a few smaller combinations that actually rank higher, but none is consistently better than the full ensemble.
Inter-annotator agreement between LLMs
To obtain additional insights into how the individual models of the ensemble diverge in their overall score predictions, we compute several measures of inter-annotator agreement. The results presented in Appendix I indicate only low to moderate agreement for most of the datasets, especially on exact overall scores, which suggests a sufficient diversity for combining outputs of these models into an ensemble.
Error analysis aggregation
The consolidator model aggregates error span annotations from multiple LLMs, which could potentially lead to an overall larger number of detected errors. To estimate the extent of over-annotation by both the ensemble and its components, we analyze the number of detected errors for output-aspect pairs rated with maximum score by human annotators. The results presented in Appendix K indeed show some tendency of the ensemble to over-annotate due to error accumulation from the individual models. However, most spans annotated by the ensemble were marked as correct in the experiment in Section 7.2. This could indicate that OpeNLGaugeens finds subtle errors which human annotators overlook, but further analysis of this discrepancy is needed.
Score aggregation
Additionally, we compare different methods of aggregating overall scores of individual LLMs to a final score in Appendix J. These results indicate that despite its simplicity, simple averaging is the most effective approach, generally providing the highest correlations with human scores.